It requires two sticks of butter to make a batch of 20 cookies. How much butter will it take to make 150
cookies?
1
Expert's answer
2020-12-17T07:34:54-0500

If it requires two sticks of butter to make 20 cookies, one can see that you need 2/20 = 0.1 (one tenth) of the stick of butter to make 1 cookie. Therefore, in order to make 150 cookies, you will need 0.1*150 = 15 sticks of butter.